Successfully communicated with the Moku using the Python API on the lab computer (Oppy 1) and my personal machine (physically in the lab connected to the lab network.)
I moved Moku Pro #01 from the closet to the rack on the floor to the left of the computer and connected it via ethernet to the switch on this rack.
I followed the instructions here: Getting Started With Python. This setup is quite straightforward, the only hiccup I encountered is that I was unable to use the command 'moku list' to find the Moku's IP address. This command requires mokucli, ('command line interface') which is a separate download found here: utilities. After installing the cli on my personal machine, the command 'moku list' results in an error 'PermissionError: [WinError 5] Access is denied,' even when running from a terminal with administrator priviliges. However, this command is not necessary because the IP address of the Moku can be aquired from one of the Ipads.
I saved a script titled 'hello_moku.py' locally on the cds virtual machine which performs a minimal communication by fetching an oscilloscope timeseries.
I made and cleaned a "U Holder" to independently individual bowtie subassemblies like the piezo assembly. This U Holder has a counterbored hole for a #8 screw to go into a post or pedestal. Screwing into a 1 inch or 1.5 inch diameter pedestal requires a 0.5 in diameter spacer, which I have ordered from Newport. The height from the mounting point of the U holder to the center of the mirror is 1 inch.
After cleaning, I did a fit check, and the piezo assembly base screws into the U Holder.
Attached is the model I used to make this part. I did not add the vent grooves. The V2 longer version I ordered from 3D Hubs is better as it is compatible with more mounts. In addition, this part should be steel so it does not flex as much when there is weight on one side.
I 3D printed a holder to hold the fiber for the FPC562 fiber paddle controller to change the polarization of the light in the fiber.
Attached is the part file and the left and right halves of the part since the single part wouldn't fit on the 3D printer bed. The top cylindrical part might need some adjustment based on the 3D printer.
So that optics aren't damaged during assembly, I built an area designated for putting together optics and mounts. The area is covered in teflon to reduce damage if a drop does occur and has a 3D printed PLA ring to keep objects inside the square. Especially for BS cubes, the post should be forked down and the mount should be screwed into the mount. See attached photo.
Attached are the SolidWorks and STL files for the three 3D printed parts for this assembly.
Last week, I damaged the corner of one of the BS Cubes (the 10:90 R:T 1550 nm cube) in the Power Distribution during assembly. Luckily, the damaged is confined to just the corner. See attached photos. In responce, I decided to build an assembly area. See the next post.
Alex and Daniel
We checked a Noliac NAC2125-H08 for shorts and found none. We measured the capacitance to be 2.4641 uF. The expected capacitance is 2.4 uF. We also measured a resistance of 3.5082 kOhm. When we pressed on the piezo, we noticed a voltage. We flagged one of the wires for polairty. See attached photos. We also noticed that the nylon tipped set screws that hold the piezo in place provide a voltage.
I looked for the dots on the Piezo we tested, and they are on the wire we DID NOT FLAG. I think we should flip which wire is flagged and then use that convention going forward.
I flagged the lead of the piezo with the black dot, which is explicitly stated by Noliac as the positive electrode. When I press on the piezo, there is a positive voltage when measuring off this positive lead.
I 3D printed a holder for the Covesion SHG and its fibers.
Attached are the SolidWorks recreation of the SHG-WGCO-M-1550-40, a custom holder, and STL files. The holder was too large to fit onto the bed of the 3D printer, so I printed it in halves.
I looed at the crusher mirror to see if there was any damage. The backside had a ring with the outline of the nub, but the front did not have any damage I could see. The plate was also indented with the AJS Screw. See attached photos.
I tested the voltage across a Noliac 2023-H08-A01 and Thorlabs PA44M3KW. When measuring the voltage accross the positive lead, there is a positive voltage under compression. For Noliac piezos, the positive lead has a black dot next to it. We should flag this positive lead as this one is. See attached photos.
.jpeg files are now image files and display. I may have accidentally removed previous .jpeg files.
I added back all my .jpeg files, and most other image files are .png
I cleaned the exterior of the following parts for the Laser Filter Cavity (LFC) using isopropanol and Kim wipes. Some tape and tape residue remains, so a more aggressive cleaning agent might be required before we bring the parts into B102 to clean the insides. See attached photos for the parts I cleaned.
Parts cleaned for LFC:
8" to 6" reducing Tee, 12" Long
6" Gate Valve (quantity 2)
End Cube and End Cube Base (quantity 2)
8" diameter bellows, 5" long (probably needs more clenaing before we bring it to B102)
Long 10" to 8" Reducer "flange"
Short 10" to 8" Reducer flange
Misc 10" Flanges
I also cleaned the following parts attached to the above parts that are not needed for the LFC:
An additional 8" to 6" reducing Tee, 12" Long
6" diameter flange, 12" long tube (quantity 2)
Following the simulation in (https://mccullerlab.com/logs/lab/index.php?callRep=11455), I decided to use an optic with a constant thickness and material, instead of the substrate having the same thickness. Thus, the whole optic (coating+substrate+coating) is 2 mm thick. For the thick coating, each layer is 100 um. For the thin coating, each layer is 20 um. Interestingly, the location of the peaks changes even though the entire optic is the same thickness. This suggests the loss angle affects the eigenfrequencies, at least in the COMSOL model.
Having any BNC plugged into the DC modulation port of the thorlabs laser turnkey raising a forest of 840 Hz lines and its multiples in the noise spectra. Lee suggests it a grounding error in the input port and can be fixed with instrumentation amplifier between the signal and the port.
[Daniel, Torrey]
Tried locking the filter cavity with and without the high voltage amplifier. This can be done by scanning on the laser frequency instead of the piezo, while still locking with piezo. From 11454 you can see the 3db roll off point is ~10 kHz. IMG_0068.png shows that the amplitudes to not differ until this point, at which point the cavity lock with the HV amplifier in place rolls off in amplitude much faster. Additionally the phase is much better with no HV amplifier.
I repeated the simulation from https://mccullerlab.com/logs/lab/index.php?callRep=11445 but with a coating thickness of 40 um instead of 20 um. I doubled the amount of mesh elements in the coating so that the mesh density of the coating remained the same. The PSD in the coating at 12.5 MHz and 13 MHz, which are roughly centered between the peaks, are 2.3x higher and 3.38x higher, respectivly. The eigenfrequency of the bulk longitudinal mode is lower since the optic is 1% thicker. For the simplest test of how the total loss angle gets calculated, I am going to go back to a low Q silicon coating.
Quick data dump in case this is needed in the future.
See attached photos for some photos of the Laser Filter Cavity SolidWorks Assembly
[Daniel, Torrey]
We tested an alternate way to hold the piezo in place for the cavity. We glued a mirror to a piezo to a 1/4" thick, 1" diameter spacer (the spacer is required to hold the assembly, the thorlabs piezo is too small). The spacer is then held in place by 3 set screws with nylon tips. See pics below. After realigning the cavity we noticed the quality of the lock is arguably worse (the alignment was much farther off compared to just replacing the piezo). Additionally, the 3.3 kHz resonance is more pronounced.
In the previous set up, where the SM1 ring and viton o-ring are used to tighten, we found some metal shavings on the viton o-ring. Potentially concerning down the road.